### v1

# Non-degenerate mixed maps and contact structures

Abstract.
We study the geometry and topology of real analytic maps $\mathbb{C}^{n}\to\mathbb{C}^{k}$, where $n>k$, regarded as mixed maps, defined below. Firstly, we give two natural families of mixed isolated complete intersection singularities, called mixed ICIS, which are interesting on their own. We consider the notion of (partial) non-degeneracy for mixed maps; we prove that these define mixed ICIS and that, under suitable conditions, admit a local Milnor fibration. Then, building on previous constructions due to Oka, we obtain natural contact structures and adapted open books on a particular class of mixed links. Finally, we look at mixed links that are diffeomorphic to holomorphic ones, and we address the problem of comparing different contact structures.
Keywords. Mixed ICIS, Real singularities, Contact structures, Open book decompositions.
Mathematics Subject Classification. 32C05, 32S55, 53D35.
The first named author was supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. The second named author had support from Mexico’s PAPIIT-UNAM proyect IN101424.
Introduction
Mixed maps are real analytic maps in complex variables and their conjugates. Perhaps the first time that such maps appeared in singularity theory was in the 1973 paper [1] by N. A’Campo where he constructed non-trivial examples of real analytic maps into $\mathbb{R}^{2}$ with a local Milnor fibration; this answered a question raised by J. Milnor in his classical book [21]. Some 25 years later, the articles [34] and [33], and later Seade’s book [35], opened a line of research on Milnor fibrations for real singularities, based on the use of mixed functions (though the name mixed was not used), and highlightenning the use of complex geometry in this setting. It was M. Oka in [25] who actually coined the name “mixed singularities” and in a series of remarkable papers showed that these inherit several properties from the holomorphic context and are objects of great interest from the topological point of view. Several authors have contributed to making this a rich and interesting theory, see for instance [8], [18], [25], [26], [27], [12], [31], to cite a few. For a general account, we refer the reader to [29].
In this work we study mixed isolated complete intersection singularities, mixed ICIS for short. We give in the first part two constructions of such maps. The first of these we call Siegel maps, as these spring by considering the space of Siegel leaves of a generic linear action of $\mathbb{C}^{k}$ in $\mathbb{C}^{n}$ in the Siegel domain. This is interesting on its own. The regular part of the mixed ICIS has a canonical complex structure induced by the linear action, and it admits a $\mathbb{C}^{*}$-action with compact quotient. This gives rise to an important class of complex manifolds known as LVM-manifolds, a type of moment-angle manifolds of importance in algebraic topology and mathematical physics.
We next look at the notion of (strong) non-degeneracy and partial non-degeneracy for mixed maps. This extends the classical notion introduced by Khovanskii in [16] for holomorphic maps, and the partial non-degeneracy for functions considered by Mondal in [22] also in the holomorphic setting, and by Bode and Quiceno in [3] for mixed functions in two variables. We determine some general properties and prove that partial non-degeneracy implies the mixed ICIS property; we also show that under additional conditions, strong non-degeneracy yields the existence of a Milnor fibration on the sphere. This gives a second class of examples provided by a construction introduced by Oka in the case of mixed hypersurfaces called mixed coverings. As a byproduct, we give the mixed version of Hamm’s complete intersections [14].
On the other hand, contact structures appear in singularity theory after the work [37] of Varchenko, who proved that links of complex analytic varieties with an isolated singularity are contact submanifolds of the sphere endowed with a canonical structure. Additionally, using the spherical Milnor fibrations, Giroux proved in [11] that they are closely related to open book decompositions. Since then, Oka and others authors have studied contact structures associated to real and complex singularities.
We extend previous constructions due to Oka in [28], giving sufficient conditions on mixed links associated with certain classes of non-degenerate mixed maps, to admit a canonical contact structure. In particular, we generalize Oka’s results, in which pullbacks of partially non-degenerate holomorphic maps are used to construct mixed maps through homogeneous mixed coverings. We show that the associated links are also contact submanifolds of the sphere, and we prove the existence of adapted open books, based on the works [6] and [5]. The argument considers a suitable change of the defining contact form developed in [6] and [28], and a formulation of a condition associated with non-isolated complex singularities in [6]. Furthermore, it happens that some classes of links defined by mixed functions are diffeomorphic to holomorphic ones. In the last part, we discuss the relation between the natural contact structure of the mixed link (when this exists) and the one induced from the holomorphic case.
1. Mixed ICIS
1.1. Definitions
A mixed map is a complex vector-valued map $F:\mathbb{C}^{n}\longrightarrow\mathbb{C}^{k}$ which is real analytic in the variables $z=(z_{1},\dots,z_{n})$ and $\mkern 3.0mu\overline{\mkern-3.0muz}=({}\mkern 3.0mu\overline{\mkern-3.0muz}_{1},\dots,{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n})\in\mathbb{C}^{n}$. If $k=1$, we denote it by $f$ and call it a mixed function. In particular, a mixed function has a series expansion of the form $\sum_{\mu,\nu}\lambda_{\mu,\nu}z^{\mu}{}\mkern 3.0mu\overline{\mkern-3.0muz}^{\nu}$. Mixed maps are those for which the coordinate functions are mixed. In this case, we have the following associated differentials:
$\partial f=\sum_{j=1}^{n}\frac{\partial f}{\partial z_{j}}dz_{j}\quad,\;\;\;{}\mkern 3.0mu\overline{\mkern-3.0mu\partial}f=\sum_{j=1}^{n}\frac{\partial f}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}\quad,\;\;\;df=\partial f+{}\mkern 3.0mu\overline{\mkern-3.0mu\partial}f.$
We define the following complex gradients:
(1.1) $Df(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\left(\frac{\partial f}{\partial z_{1}},\dots,\frac{\partial f}{\partial z_{n}}\right),\;{}\mkern 3.0mu\overline{\mkern-3.0muD}f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\left(\frac{\partial f}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}},\dots,\frac{\partial f}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n}}\right)\in\mathbb{C}^{n}.$
A critical point of a mixed map germ is a point for which the rank of the Jacobian matrix is not maximal and we call it a mixed singularity. The set of critical points of a map $F$ is denoted by $\Sigma_{F}$ and its zero set $F^{-1}(0)$ by $V_{F}$. We have the following characterization of mixed singularities stated in [31, Proposition 4].
Proposition 1.1.
Let $F=(f^{1},\dots,f^{k}):\mathbb{C}^{n}\longrightarrow\mathbb{C}^{k}$ be a mixed map germ. Then $a\in\mathbb{C}^{n}$ is a mixed singularity if and only if there exists $\alpha_{1},\dots,\alpha_{k}\in\mathbb{C}$ non-simultaneously vanishing such that
(1.2) $\alpha_{1}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{1}(a)+\dots+\alpha_{k}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{k}(a)={}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}{}\mkern 3.0mu\overline{\mkern-3.0muDf_{1}}(a)+\dots+{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}{}\mkern 3.0mu\overline{\mkern-3.0muDf_{k}}(a).$
This generalizes [25, Proposition 1], which states the following.
Corollary 1.2.
Let $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ be a mixed function germ. Then $a\in\mathbb{C}^{n}$ is a mixed singularity if and only if there exists a complex number $\alpha$ such that $\|\alpha\|=1$ and $\mkern 3.0mu\overline{\mkern-3.0muDf}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})=\alpha{}\mkern 3.0mu\overline{\mkern-3.0muD}f(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})$.
Definition 1.3.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map such that $n>k$ and $V_{F}$ has positive dimension. We say that $F$ is a mixed isolated complete intersection singularity, mixed ICIS for short, if $\Sigma_{F}\cap V_{F}=\{0\}$.
Remark 1.4.
Under the condition above, the map $F$ is regular at every point $p\in V_{F}\setminus\{0\}$. Thus, $V_{F}$ has the correct dimension nkn-k, or equivalently, it is a geometric complete intersection. In addition, observe that if $F$ is a holomorphic map, it coincides with the geometric characterization of isolated complete intersection singularities.
Next, we describe two constructions of mixed ICIS.
1.2. Linear actions on $\mathbb{C}^{n}$
Our first construction of mixed ICIS springs from complex geometry and dynamics. Recall that a linear vector field $F$,
$\mathbb{C}^{n}\ni\left(z_{1},\cdots,z_{n}\right)\stackrel{{\scriptstyle F}}{{\longmapsto}}\sum_{i=1}^{n}\lambda_{i}z_{i}\frac{\partial}{\partial z_{i}}$
in $\mathbb{C}^{n}$ is in the Siegel domain if the eigenvalues $\lambda_{i}$ are complex numbers such that their convex hull $\mathcal{H}\left(\Lambda_{1},\cdots,\Lambda_{n}\right)$ contains the origin $0\in\mathbb{C}^{n}$. It defines a holomorphic linear action of $\mathbb{C}$ in $\mathbb{C}^{n}$ with 0 as the unique fixed point. The orbits define a 1-dimensional holomorphic foliation $\mathcal{F}$, singular at 0. Let $\mathcal{T}_{F}$ be the set of points where the leaves of $\mathcal{F}$ are tangent to the foliation given by all (2n1)(2n-1)-spheres in $\mathbb{C}^{n}\setminus\{0\}$ centered at 0. It is clear that a point $z=\left(z_{1},\ldots,z_{n}\right)\in\mathbb{C}^{n}-\{0\}$ is in $\mathcal{T}_{F}$ if and only if the Hermitian product $\langle F(z),z\rangle$ vanishes. That is:
$\sum_{i=1}^{n}\lambda_{i}z_{i}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}=0,$
or equivalently where the real and the imaginary parts vanish:
$\sum_{i=1}^{n}\operatorname{Re}\left(\lambda_{i}\right)\left|z_{i}\right|^{2}=0\quad\text{ and }\quad\sum_{i=1}^{n}\operatorname{Im}\left(\lambda_{i}\right)\left|z_{i}\right|^{2}=0.$
Then $\psi(z)=\langle F(z),z\rangle$ is a mixed function and if we assume further the following genericity condition:
$i\neq j\Rightarrow\lambda_{i}\notin\mathbb{R}\lambda_{j},\text{ for all }i,j=1,\ldots,n,$
which can only happen for $n\geq 3$, then we know from [4] that there is an open dense set in $\mathbb{C}^{n}\setminus\{0\}$ of Siegel leaves, each such leaf being a copy of $\mathbb{C}$ embedded in $\mathbb{C}^{n}$ with a unique point if $\mathcal{T}_{F}$, which is the point in its leaf of minimal distance to 0. Moreover, $\mathcal{T}_{F}\setminus\{0\}$ is a (2n2)(2n-2)-dimensional smooth manifold that parameterizes the space of Siegel leaves. As noticed in [18], $\mathcal{T}_{F}\setminus\{0\}$ transversal everywhere to the leaves of $\mathcal{T}_{F}$ and therefore it inherits a canonical holomorphic structure from that of $\mathcal{F}$.
More generally, consider a linear action $\mathcal{A}$ of $\mathbb{C}^{k}$ on $\mathbb{C}^{n}$, where $0<2k<n$, generated by $k$ holomorphic linear commuting vector fields.
$\left(z_{1},\cdots,z_{n}\right)\stackrel{{\scriptstyle F^{j}}}{{\longmapsto}}\sum_{i=1}^{n}\lambda_{ij}z_{i}\frac{\partial}{\partial z_{i}}\quad,\quad j=1,\ldots,k\,.$
Let us assume the genericity condition that the matrix $M=\left(\lambda_{ij}\right)$, with $i\in\{1,\cdots,n\}$ and $j\in\{1,\cdots,k\}$, has rank $k$. Let $\mathcal{F}$ be the complex foliation on $\mathbb{C}^{n}$ whose leaves are the orbits of this action, and let $\Lambda:=\left(\Lambda_{1},\cdots,\Lambda_{n}\right)$ be the $n$-tuple of vectors in $\mathbb{C}^{k}$ defined by $\Lambda_{i}=\left(\lambda_{i1},\cdots,\lambda_{ik}\right)$ for $i=1,\cdots,n$. Following [19], we define:
Definition 1.5.
  1. (1)
    The action is in the Siegel domain if the convex hull of $\left(\Lambda_{1},\cdots,\Lambda_{n}\right)$ in $\mathbb{C}^{k}$ contains the origin:
    $0\in\mathcal{H}\left(\Lambda_{1},\cdots,\Lambda_{n}\right).$
  2. (2)
    It is admissible if it is in the Siegel domain and satisfies the following weak hyperbolicity condition: For every $2k$-tuple of integers $i_{1},\ldots,i_{2k}$ such that $1\leq i_{1}<\ldots<i_{2k}\leq n$, we have that $0\notin\mathcal{H}\left(\Lambda_{i_{1}},\ldots,\Lambda_{i_{2k}}\right)$. In this case we say that the $k$-frame $\mathfrak{F}:=\left(F^{1},\ldots,F^{k}\right)$ of commuting linear vector fields is admissible.
[…]
$\left\langle F^{j}(z),z\right\rangle:=\sum_{i=1}^{n}\lambda_{i}^{j}\left|z_{i}\right|^{2}=0\quad,\quad\forall j=1,\ldots,k\,.$
Notice that each of these is a mixed function $\psi_{j}$ and we know from [19] that $\mathcal{T}^{*}$ is smooth of real codimension $2k$. Hence the variety $V:=V_{\mathfrak{F}}$ defined by $\left(\psi_{1},\ldots,\psi_{j}\right)$ is a mixed ICIS. Moreover, $\mathcal{T}_{\mathfrak{F}}^{*}$ is everywhere transversal to the leaves of $\mathcal{F}$ so, by [13], $\mathcal{T}_{\mathfrak{F}}^{*}$ has a holomorphic structure inherited from that in $\mathcal{F}$. This does not mean that $\mathcal{T}_{\mathfrak{F}}^{*}$ is a complex submanifold of $\mathbb{C}^{n}$, neither that $\mathcal{T}_{\mathfrak{F}}$ is a complex singularity.
We summarize this discussion in the following theorem. This extends to complete intersections the method from [34] and [33] to construct real analytic singularities via complex geometry, and it is a reformulation of the results in [4] and [18] for $k=1$ and from [19] for $k>1$.
Theorem 1.6.
Let $\mathfrak{F}:=\left(F^{1},\ldots,F^{k}\right)$ be an admissible frame of $k$ commuting linear vector fields in the Siegel domain. Define $k$ mixed functions $\mathbb{C}^{n}\rightarrow\mathbb{C}$ by:
$\psi^{j}(z)=\left\langle F^{j}(z),z\right\rangle:=\sum_{i=1}^{n}\lambda_{ij}\left|z_{i}\right|^{2}.$
Then:
  1. (1)
    The map $\Psi_{\mathfrak{F}}=\left(\psi^{1},\ldots,\psi^{k}\right)$ is a mixed map and $\mathcal{T}_{\mathfrak{F}}=\Psi_{\mathfrak{F}}^{-1}(0,\ldots,0)$ is a mixed ICIS.
  2. (2)
[…]
Whence, we define:
Definition 1.7.
Let $\mathfrak{F}:=\left(F^{1},\ldots,F^{k}\right)$ be an admissible frame of $k$ commuting linear vector fields in the Siegel domain, and let $\Psi_{\mathfrak{F}}=\left(\psi^{1},\ldots,\psi^{k}\right)$ be as above. We call $\Psi_{\mathfrak{F}}$ a Siegel complete intersection map.
Remark 1.8.
We know from [18, 19] that the variety $\mathcal{T}_{\mathfrak{F}}$ admits a canonical $\mathbb{C}^{*}$-action, which is a polar action in the sense of [8] and it preserves the complex structure in $\mathcal{T}_{\mathfrak{F}}^{*}$. The quotient $\mathcal{T}_{\mathfrak{F}}^{*}/\mathbb{C}^{*}$ is a compact complex orbifold with a very interesting geometry and topology. These give rise to the LVM-manifolds, a special type of moment-angle manifolds. We refer to [20] for a thorough account on the subject.
Remark 1.9.
Notice that one may consider a vector field $F=\left(\lambda_{1}z_{\sigma(1)}^{a_{\sigma(1)}},\ldots,\lambda_{n}z_{\sigma(n)}^{a_{\sigma(n)}}\right)$, where $\sigma$ is a permutation of $\{1,\dots,n\}$, and the real analytic function $z\mapsto\langle F(z),z\rangle$. The zero set of this function describes the points where $F$ is tangent to the spheres centered at 0. Re-labeling the variables and assuming $\lambda_{i}=1$ for simplicity, this takes the form:
(1.3) $\Psi_{F}=z_{1}^{a_{1}}\bar{z}_{\sigma(1)}+\cdots+z_{n}^{a_{n}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{\sigma(n)}.$
These are the twisted Pham-Brieskorn polynomials from [34] and [33], where it is proved that if the $a_{i}$ are all $\geq 2$, then these have a unique critical point at 0 and they have a Milnor fibration. This was the birth of the theory of mixed functions. In fact these singularities have a canonical action of $\mathbb{R}\times\mathbb{S}^{1}$, later called a polar action [8]. When the permutation $\sigma$ is the identity, these were called mixed Pham-Brieskorn polynomials in [25]. Of course one may consider now several of these equations and ask under which conditions the resulting map is a mixed complete intersection. A particular case corresponds to the mixed Hamm complete intersections, discussed in Subsection 2.2.
1.3. Mixed coverings
We now introduce a method due to Oka in [28] that allows constructing mixed maps out from holomorphic ones. He used this to construct interesting mixed functions and mixed hypersurfaces. This method also works for complete intersections.
Let $\textbf{a}=(a_{1},\dots,a_{n})$ and $\textbf{b}=(b_{1},\dots,b_{n})$ be vectors of positive integers such that $a_{i}\neq b_{i}$ for all $i=1,\dots,n$. A mixed covering $\phi_{\textbf{a},\textbf{b}}$ is the map germ $\phi_{\textbf{a},\textbf{b}}:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ defined by
$\phi_{\textbf{a},\textbf{b}}(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\left(w_{1}^{a_{1}}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{1}^{b_{1}},\dots,w_{n}^{a_{n}}{}\mkern 3.0mu\overline{\mkern-3.0muw}^{b_{n}}\right).$
If there exist positive integers $a\neq b$ such that $a_{i}=a$ and $b_{i}=b$ for all $i$, then $\phi_{\textbf{a},\textbf{b}}$ is called a homogeneous mixed covering and denoted by $\phi_{a,b}$. Observe that $\phi:(\mathbb{C}^{*n},0)\longrightarrow(\mathbb{C}^{*n},0)$ is a diffeomorphism. Notice yet the similarity of this construction with the mixed maps (1.3).
1.4. Algebraic ICIS
For the results stated in this subsection, we refer the reader to [38] and [32]. Recall that a holomorphic map germ $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ is an ICIS if and only if it is $C^{\infty}-\mathcal{K}$-finitely determined, where $\mathcal{K}$ is the contact group of Mather. On the other hand, if $F$ is real analytic, then it defines a mixed ICIS if and only if it is $C^{l}-\mathcal{K}$-finitely determined for all $l\in[0,\infty)$. As we shall see, there exist mixed ICIS which are not $C^{\infty}-\mathcal{K}$-finitely determined. This leads to the following definition.
Definition 1.10.
We say that a mixed map germ $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ is an algebraic ICIS if it is $C^{\infty}-\mathcal{K}$-finitely determined.
This class encompasses the holomorphic ICIS and real analytic maps $(\mathbb{R}^{2n},0)\longrightarrow(\mathbb{R}^{2k},0)$ that are $C^{\infty}-\mathcal{K}$-finitely determined. Moreover, by the characterization mentioned above, an algebraic ICIS is a mixed ICIS as in Definition 1.3. We shall see that this is not a general property of the mixed setting.
Example 1.11.
Let $\Psi$ be a Siegel complete intersection map determined by an admissible configuration $\Lambda=(\Lambda_{1},\dots,\Lambda_{n})$. We claim that it is not an algebraic ICIS. For this, consider $\Psi:(\mathbb{R}^{2n},0)\longrightarrow(\mathbb{R}^{2k},0)$ and its complexification $\Psi_{\mathbb{C}}:(\mathbb{C}^{2n},0)\longrightarrow(\mathbb{C}^{2k},0)$. Its coordinate functions are:
$\operatorname{Re}\Psi^{i}_{\mathbb{C}}=\sum_{j=1}^{n}\operatorname{Re}\lambda_{ij}\left(\xi_{i}^{1}+\xi_{i}^{2}\right)\quad\hbox{and}\quad\operatorname{Im}\Psi^{i}_{\mathbb{C}}=\sum_{j=1}^{n}\operatorname{Im}\lambda_{ij}\left(\xi_{i}^{1}+\xi_{i}^{2}\right)\,$,
where $\xi_{i}^{1},\xi_{i}^{2}$ are complex variables. Take the union $L$ of the complex lines $L_{i}=\{\xi_{i}^{1}=\pm{\bf i}\xi_{i}^{2}\}$. At a point $p\in L$, $\Psi_{\mathbb{C}}(p)=0$ and its Jacobian matrix has pairwise linearly dependent columns. In other words, $L\subset\Psi_{\mathbb{C}}^{-1}(0)\cap\Sigma_{\Psi_{\mathbb{C}}}$ and $\Psi$ is not an algebraic ICIS because this property does not hold for its complexification.
Definition 1.12.
  1. (1)
    A mixed monomial $f_{\mu,\nu}=\lambda_{\mu,\nu}z^{\mu}{}\mkern 3.0mu\overline{\mkern-3.0muz}^{\nu}$ is called purely mixed with respect to $z_{i}$ if $\mu_{i},\nu_{i}\geq 1$ and $\mu_{i}+\nu_{j}\geq 3$.
  2. (2)
    A mixed function $f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{\mu,\nu}f_{\mu,\nu}$ is called purely mixed if every mononomial $f_{\mu,\nu}$ is purely mixed with respect to some variable $z_{i}$.
Proposition 1.13.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map such that $f^{i}$ is purely mixed for every $i=1,\dots,k$. Then $F$ is not an algebraic mixed ICIS.
Proof.
The proof essentially follows by the same argument as in Example 1.11. Let $f$ be a purely mixed coordinate function. A straightforward computation shows that for each monomial $f_{\mu,\nu}$, the partial derivatives have the following form:
$\frac{\partial}{\partial z_{i}}\left(2\operatorname{Re}f_{\mu,\nu}\right)=\frac{\partial}{\partial z_{i}}\left(f_{\mu,\nu}+{}\mkern 3.0mu\overline{\mkern-3.0muf_{\mu,\nu}}\right)=\|z_{j}\|^{2}h_{\mu,\nu},$
for some index $j$ and a mixed function $h_{\mu,\nu}$. Similar expressions hold for $\operatorname{Im}f_{\mu,\nu}$ and the derivatives with respect to $\mkern 3.0mu\overline{\mkern-3.0muz}_{i}$. Applying it to every monomial, one can see that $f_{\mathbb{C}}$ and its derivative vanish on the set $L$ consisting of the union of the zero sets correspondent to the complexification of $\|z_{i}\|^{2}$. Therefore, since each coordinate function $f^{i}$ of $F$ is purely mixed, $L\subset F_{\mathbb{C}}^{-1}(0)\cap\Sigma_{F_{\mathbb{C}}}$ and the result follows. ∎
As a consequence, mixed maps constructed as the pullback of mixed coverings are never algebraic mixed ICIS. This illustrates the substantial differences between the mixed and holomorphic settings (see Subsection 3.2).
2. Non-degeneracy
In this section we extend the notions of non-degeneracy for mixed maps. A mixed function germ is denoted by $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ and a mixed map germ by $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$. From now on we assume that $V_{f}$ and $V_{F}$ have positive dimension in $\mathbb{C}^{n}$.
In [26], M. Oka introduced the notion of the Newton polyhedra of a mixed function germ $f$, denoted by $\Gamma^{+}(f)$, and defined the condition of (strong) non-degeneracy. If $f$ is holomorphic, these two notions coincide with the classical definitions due to Kushnirenko in [17].
Let us fix some notations. Let $I\subset\{1,\dots,n\}$ be a non-empty subset. We define $\mathbb{C}^{I}=\{z\in\mathbb{C}^{n}:z_{i}=0\;\forall\;i\notin I\}$. We denote the face function of $f$ with respect to a vector $P$ of positive integers as $f_{P}$ and its restriction to $\mathbb{C}^{I}$ by $f^{I}$. A Newton polyhedra $\Gamma^{+}$ is called convenient if it intersects each non-empty subspace $\mathbb{R}^{I}$. P. Mondal in [22] introduced the following condition, whose main advantage is that it avoids the convenience hypothesis on the Newton polyhedron in some applications.
Definition 2.1.
A holomorphic function $f(z)$ is called partially non-degenerate if for every non-empty subset $I\subset\{1,\dots,n\}$ and every vector $P$ of positive integers, the vector field $(Df)^{I}_{P}(a)\neq 0$ for all $a\in\mathbb{C}^{*n}$.
For holomorphic function germs $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ these notions are related as follows. If $f$ is non-degenerate and convenient, then $f$ is partially non-degenerate by [22, Proposition X1.7]. Moreover, if $f$ is partially non-degenerate, then it has an isolated singularity at the origin as proved in [22, Theorem X1.3]. However, there are examples in which $f$ is degenerate but partially non-degenerate, see [22, Example XI.6] and Example 2.8. More generally, these implications depend on the fact that $(Df)^{I}_{P}$ and $(Df_{P})^{I}$ do not coincide in general for a subset $I\subset\{1,\dots,n\}$ and $P$ a vector of positive integers.
Lemma 2.2.
Let $f(z)$ be a partially non-degenerate holomorphic function. There does not exist a nonzero real analytic curve $w(t):(0,1]\longrightarrow\mathbb{C}^{n}\setminus\{0\}$ such that $\lim_{t\to 0}w(t)=0$ and $f_{z_{j}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)\equiv 0$ for all $j=1,\dots,n$.
Proof.
Suppose the existence of such a curve. First, define $I=\{j:w_{j}(t)\not\equiv 0\}$. For each $j\in I$, we write the coordinate $w_{j}$ of $w$ as $w_{j}(t)=a_{j}t^{p_{j}}+o(t)$, where $a_{j},p_{j}\neq 0$ and $o(t)$ denotes higher order terms. One can see that
$\frac{\partial f}{\partial z_{j}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)=\left(\frac{\partial f}{\partial z_{j}}\right)^{I}_{P}(a)t^{d-p_{j}}+o(t),$
where $P=(p_{1},\dots,p_{n})$ and $d$ is the weighted homogeneous degree of $f_{P}$. We conclude that $\left(Df\right)_{P}^{I}(a)=0$, which contradicts the partial non-degeneracy condition. ∎
Mixed functions with an isolated singularity at the origin can be constructed from holomorphic ones as follows.
Proposition 2.3.
Let $f(z)$ be a partially non-degenerate holomorphic function and $g=\phi_{\textbf{a},\textbf{b}}^{*}f$ for some mixed covering $\phi_{\textbf{a},\textbf{b}}(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$. Then $g(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$ has an isolated mixed singularity at the origin.
Proof.
Otherwise, by the Curve Selection Lemma and Corollary 1.2, one can find real analytic curves $\lambda(t)\subset\mathbb{S}^{1}$ and $w(t)\subset\Sigma_{g}$ such that $\lim_{t\to}w(t)=0$ and
$\overline{f_{z_{j}}}(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))a_{j}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}(t)^{a_{j}-1}w_{j}^{b_{j}}=\lambda(t)b_{j}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{b_{j}-1}w_{j}(t)^{a_{j}}f_{z_{j}}\left(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\right).$
Let $I=\{j:w_{j}(t)\not\equiv 0\}$. Since $a_{j}\neq b_{j}$ and taking norms, the expression above implies that $f_{z_{j}}$ vanishes on $\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$ for all $j\in I$. This leads to a contradiction with Lemma 2.2. ∎
2.1. Mixed maps
We extend the definition of non-degeneracy of holomorphic maps in [16] to the case of mixed map germs. Similar notions are considered, for example, in [23], [2], [24], and at infinity by [7] and [36].
Definition 2.4.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map. We say that $F$ is non-degenerate with respect to the Newton boundaries $\Gamma(f^{1}),\dots,\Gamma(f^{k})$ if, for every vector $P$ of positive integers, the following condition is verified: at each point $p\in\mathbb{C}^{*n}$ such that $F_{P}(p)=0$, the differentials $Df_{P}^{1},{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{1},\dots,Df_{P}^{k},{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{k}$ do not satisfy a relation as in (1.2). In addition, we say that $F$ is strongly non-degenerate if the previous condition holds for any point $p\in\mathbb{C}^{*n}$.
Example 2.5.
Recall the Siegel complete intersection map $\Psi_{\mathcal{F}}$ in Definition 1.7. Let $\Lambda=(\Lambda_{1},\dots,\Lambda_{n})$ the $n$-tuple of $\mathbb{C}^{k}$-vector defining this mixed ICIS. Let us suppose that for all $m$-tuples $(i_{1},\dots,i_{m})$ of integers in $\{1,\dots,n\}$, where $m\geq 2k$, the set $(\Lambda_{i_{1}},\dots,\Lambda_{i_{m}})$ is an admissible configuration. Then the mixed map $\Psi_{\mathcal{F}}$ becomes non-degenerate, since its face functions are the restrictions to subspaces $\mathbb{C}^{I}$ and $\Psi_{\mathcal{F}}^{I}=0$ has no solution in $\mathbb{C}^{*I}$ if $|I|<k$.
Non-degeneracy notions are well known for holomorphic maps and mixed functions and form a generic class. We indicate a simple procedure to construct new maps with the same properties.
Example 2.6.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ and $G:(\mathbb{C}^{m},0)\longrightarrow(\mathbb{C}^{l},0)$ be (strongly) non-degenerate mixed maps and consider the map $H=\left(F,G\right):(\mathbb{C}^{n+m},0)\longrightarrow(\mathbb{C}^{k+l},0)$ formed by $F$ and $G$ on separable variables. Since the derivative of the face map $H_{P}$ has a diagonal form, it has maximal rank if and only if both derivatives of $F_{P}$ and $G_{P}$ have maximal rank, for every vector of positive integers $P$.
We extend partial non-degeneracy for mixed maps. We remark that this definition for mixed functions on two variables was introduced in [3].
Definition 2.7.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map. We say that $F$ is partially non-degenerate with respect to the Newton boundaries $\Gamma(f^{1}),\dots,\Gamma(f^{k})$ if, for every vector $P$ of positive integers and every nonempty subset $I\subset\{1,\dots,n\}$, the following condition is verified: at each point $p\in\mathbb{C}^{*n}$ such that $F_{P}^{I}(p)=0$, the differentials $(Df^{1})_{P}^{I}\;,\;({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1})_{P}^{I}\,,\dots,(Df^{k})_{P}^{I}\;,\;({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k})_{P}^{I}$ do not satisfy a relation as in (1.2).
Example 2.8.
Let $F:(\mathbb{C}^{3},0)\longrightarrow(\mathbb{C}^{2},0)$ be given by $F=(z_{1}+(z_{2}+z_{3})^{2},z_{1}^{2}+z_{2}^{2}+z_{3}^{2})$. The derivative $DF$ of $F$ consists of a constant and terms of order 1, so $DF=(DF)_{P}$ for every vector $P$ of positive integers. Moreover, for every non-empty $I\subsetneq\{1,2,3\}$, $(DF)^{I}_{P}$ does not have maximal rank only at points $p\in F_{P}^{-1}(0)$ in the complement of $\mathbb{C}^{*I}$. On the other hand, we may choose a vector $P$ so that the first coordinate function of $F_{P}$ is $(z_{2}+z_{3})^{2}$. In this case, $DF_{P}$ does not have maximal rank at points $p\in F_{P}^{-1}(0)\cap\mathbb{C}^{*3}$. Thus, $F$ is degenerate but partially non-degenerate.
Proposition 2.9.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ.
  1. (1)
    If $F$ is non-degenarate, then $F^{I}$ is also non-degenerate, where $I\subset\{1,\dots,n\}$ is such that $f^{i,I}\not\equiv 0$ for every $i=1,\dots,k$.
  2. (2)
    If $F$ is (strongly, partially) non-degenerate and $\phi$ is a mixed covering, then the pullback $G=\phi^{*}F$ is also (strongly, partially) non-degenerate.
  3. (3)
    If $F$ is non-degenerate and convenient, then it is partially non-degenerate.
Proof.
For the first item, we follow the proof in [26, Proposition 7] for mixed functions. Let $P$ be a vector of positive integers and denote $\left(f^{i,I}\right)_{P}=f^{i,I}_{P}$ for every $i=1,\dots,k$. Let $Q_{i}=(q^{i}_{1},\dots,q^{i}_{n})$ be a vector such that $q^{i}_{j}=p_{i}$ if $i\in I$ and $q^{i}_{j}=v^{i}_{j}$ if $i\notin I$, where $v^{i}_{j}$ are positive integers. If $v^{i}_{j}$ are sufficiently large, then
$f^{i}_{Q_{i}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\left(f^{i,I}\right)_{P}(z_{I},{}\mkern 3.0mu\overline{\mkern-3.0muz}_{I}),$
where $(z_{I},{}\mkern 3.0mu\overline{\mkern-3.0muz}_{I})=(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\cap\mathbb{C}^{I}$. We may take $v^{i}=\max_{j}\{v^{i}_{j}\}$ and define $Q=(q_{1},\dots,q_{n})$ such that $q_{i}=p_{i}$ if $i\in I$ and $q_{i}=v_{i}$ if $i\notin I$. It follows that $F^{I}_{P}=F_{Q}$ and the non-degeneracy of $F$ is translated to the non-degeneracy of $F^{I}$.
In the second item, for each vector $P$ of positive integers one can see that $G_{P}=\phi^{*}F_{P}$. Since $\phi$ is a diffeomorphism on $\mathbb{C}^{*n}$, the assertion for (strong) non-degeneracy follows. On the other hand, for every non-empty subset $I\subset\{1,\dots,n\}$, the derivatives are related by $(DF)_{P}^{I}=(DG)_{P}^{I}\circ D\phi^{I}$. Since $\phi^{I}$ is a diffeomorphism of $\mathbb{C}^{*I}$ we conclude the proof for partially non-degenerate maps.
The last assertion is a consequence of the following remark (see [22, Proposition X1.7]). Let $f$ be a coordinate function of $F$. If $\left(f_{P}^{I}\right)_{z_{i}}\not\equiv 0$, then it is equal to $\left(f_{z_{i}}\right)^{I}_{P}$, where $I\subset\{1,\dots,n\}$ is non-empty. The analogous property holds for partial derivatives with respect to $\mkern 3.0mu\overline{\mkern-3.0muz}_{i}$. By the convenience hypothesis we can conclude the assertion by the first item.
Item 2 of the proposition above is analogous to [28, Proposition 6] for mixed functions. One has that non-degeneracy property for holomorphic maps is a general condition by [16]. Therefore, this assertion provides several examples of non-degenerate mixed maps. The main result of this section relates non-degeneracy and ICIS properties.
Theorem 2.10.
Let $F=(f^{1},\dots,f^{k}):\mathbb{C}^{n}\longrightarrow\mathbb{C}^{k}$ be a partially non-degenerate mixed map such that $V_{F}$ has positive dimension. Then $F$ is a mixed ICIS.
Proof.
On the contrary, by the Curve Selection Lemma and Proposition 1.1, there exists a real analytic curve $w:(0,1]\longrightarrow(\Sigma_{F}\cap V_{F})\setminus\{0\}$ such that $\lim_{t\to 0}w(t)=0$, and real analytic curves $\alpha_{1},\dots,\alpha_{k}\subset\mathbb{C}$ non-simultaneously vanishing such that
$\alpha_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)+\alpha_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)={}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{1}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)+\dots+{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{k}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right).$
Let us suppose that $I=\{i:w_{i}(t)\not\equiv 0\}$ and without loss of generality, $\alpha_{i}(t)\not\equiv 0$ for all $i=1,\dots,k$. Consider the real analytic expansion of the curves:
$w_{i}(t)=a_{i}t^{p_{i}}+o(t),\;a_{i}\in\mathbb{C}^{*},p_{i}\geq 1,$
$\alpha_{i}(t)=c_{i}t^{q_{i}}+o(t),\;c_{i}\in\mathbb{C}^{*},q_{i}\geq 1,$
for all $i\in I$, where $o(t)$ denotes higher order degree terms. Let $d_{i}$ be the weighted homogeneous degree of the face functions $f^{i}_{P}$, for $i=1,\dots,k$. For each $i$, the following equations hold:
(2.1) $\begin{split}\alpha_{1}(t)\left(\left(\frac{\partial f^{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\right)^{I}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}+o(t)\right)+&\dots+\alpha_{k}(t)\left(\left(\frac{\partial f^{k}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\right)_{P}^{I}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right)\\ &=\\ {}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t)\left(\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}}\right)^{I}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}+o(t)\right)+&\dots+{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t)\left(\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{k}}{\partial z_{i}}}\right)^{I}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right),\end{split}$
We may suppose $q_{1}+d_{1}\leq\dots\leq q_{k}+d_{k}$. Let $J$ denote the indices $j=1,\dots,l$, with $l\leq k$, for which $q_{1}+d_{1}=\dots=q_{l}+d_{l}$. Comparing the orders of both sides, it follows that the differentials $\left({}\mkern 3.0mu\overline{\mkern-3.0muDf^{1}}\right)^{I}_{P},\dots,\left({}\mkern 3.0mu\overline{\mkern-3.0muDf^{k}}\right)^{I}_{P}$, $\left({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1}\right)^{I}_{P},\dots,\left({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k}\right)^{I}_{P}$ are linearly dependent at $a\in\left(F^{I}_{P}\right)^{-1}(0)\cap\mathbb{C}^{*m}$, where we set zero for coefficients whose indices do not belong to $J$. This leads to a contradiction with the partial non-degeneracy of $F$. ∎
2.2. Mixed Hamm Complete Intersections
A particular example of genuine mixed map defining an ICIS can be obtained from the previous constructions as follows. Let $\textbf{a}=(a_{1},\dots,a_{n}),$ be a vector of positive integers and $\Lambda=(\lambda_{ij})$ a complex matrix of order $n\times k$. For each $i=1,\dots,k$, let $f^{i}=\sum_{j=1}^{n}\lambda_{ij}z_{j}^{a_{j}}$ be a complex Pham-Brieskorn polynomial. Hamm showed in [14] that for a sufficiently general matrix, the map germ $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ defines an ICIS, which we shall refer to as Hamm ICIS.
Let $\textbf{b}=(b_{1},\dots,b_{n})$ be a second vector of non negative integers and consider the mixed Pham-Brieskorn polynomial $g^{i}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{j=1}^{n}\lambda_{ij}z_{i}^{a_{i}+b_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}^{b_{i}}$. This type of mixed functions is a particular case of the construction discussed in Subsection 1.2. If we require all $k\times k$-minors of $\Lambda$ being nonzero, the map $H:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ whose coordinate functions are $h^{i}(z)=\sum_{j=1}^{n}\lambda_{ij}z_{j}$ is non-degenerate. By Proposition 2.9 and Theorem 2.10 we conclude that the mixed Hamm map defines an ICIS, which we call mixed Hamm ICIS. Notice that $G$ is not an algebraic ICIS by Proposition 1.13.
In [33, Theorem 4.1] it is shown that complex and mixed Pham-Brieskorn polynomials are topologically equivalent, and this assertion easily applies for the maps constructed above. Furthermore, Oka proved in [27] that the links are smoothly equivalent. In this section, we extend this result for mixed Hamm ICIS. We fix the vectors a and b of integers. For each $i=1,\dots,k$, let $f^{i}$ and $g^{i}$ be mixed and complex Pham-Brieskorn polynomials, respectively, as before. Define the following family of mixed maps:
$G_{t}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=(1-t)G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})+tF(z,{}\mkern 3.0mu\overline{\mkern-3.0muz}),$
where $F=(f^{1},\dots,f^{k})$ and $G=(g^{1},\dots,g^{k})$. Let us denote $V_{t}^{i}=\left(g_{t}^{i}\right)^{-1}(0)$, where $g_{t}^{i}$ are the coordinate functions of $G_{t}$, and $V_{t}=G_{t}^{-1}(0)$. We fix the notation $\mathbb{C}^{n}_{*}:=\mathbb{C}^{n}\setminus\{0\}$.
Lemma 2.11.
Let $G_{t}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ be as above, where $0\leq t\leq 1$. The following facts hold true.
  1. (1)
    The map $G_{t}$ is a mixed ICIS.
  2. (2)
    The variety $V_{t}$ intersects the sphere $\mathbb{S}_{r}^{2n-1}$ transversely for any $r>0$.
  3. (3)
    Let $r>0$ be fixed. Then there exists a family of diffeomorphisms
    $\psi_{t}:(\mathbb{B}_{r}^{2n},E_{0}(r))\longrightarrow(\mathbb{B}_{r}^{2n},E_{t}(r)),$
    where $E_{t}(r)=\{z\in\mathbb{C}^{n}_{*}:G_{t}(z)=0,\|z\|\leq r\}$. Moreover, it restricts also as diffeomorphisms
    $\psi_{t}:(\mathbb{S}_{r}^{2n-1},\partial E_{0}(r))\longrightarrow(\mathbb{S}_{r}^{2n-1},\partial E_{t}(r)),$
    where $\partial E_{t}(r)=\{z\in\mathbb{C}^{n}_{*}:G_{t}(z)=0,\|z\|=r\}$.
Proof.
In the first item, for $t=0$ and $t=1$ the assertion is already proved. In the other cases, it is enough to notice that the Newton polyhedron of $g_{t}^{i}$ is the polyhedron of a complex Pham-Brieskorn polynomial and we reduce to a case already settled. The second item is a consequence of [27, Lemma 2], which shows the assertion for each $V_{t}^{i}$ and thus for the intersection $V_{t}=V_{t}^{1}\cap\dots\cap V_{t}^{k}$. The last statement follows from Ehresmann fibration theorem for subbundles applied to the canonical projections
$\pi:E(r)\times I\longrightarrow I\quad\text{and}\quad\partial\pi:\partial E(r)\times I\longrightarrow I,$
where
$E(r)$ $=\{(z,t)\in\mathbb{C}^{n}_{*}\times I:G_{t}(z)=0,\|z\|\leq r\},$
$\partial E(r)$ $=\{(z,t)\in\mathbb{C}^{n}_{*}\times I:G_{t}(z)=0,\|z\|=r\}.$
Let $I\subset\{1,\dots,n\}$ be a non-empty subset such that $|I|>k$. As before, let us denote with an upper index $I$ the restriction of the sets and maps to the subspace $\mathbb{C}^{I}=\{z\in\mathbb{C}^{n}:z_{i}=0\;\text{if}\;i\notin I\}$. Observe that, by non-degeneracy, the restrictions of the maps and sets in Lemma 2.11 to $\mathbb{C}^{I}$ share the same properties. Then, we may conclude the discussion as follows.
Theorem 2.12.
Let $F$ and $G$ be the complex and mixed Hamm ICIS, respectively, and let $r>0$ be fixed.
  1. (1)
    There exists a diffeomorphism
    $\psi:(\mathbb{S}_{r}^{2n-1},K_{G})\longrightarrow(\mathbb{S}_{r}^{2n-1},K_{F}),$
    where $K_{G}$ and $K_{F}$ are the links defined by $F$ and $G$, respectively.
  2. (2)
    Let $I\subset\{1,\dots,n\}$ such that $|I|>k$. Then the map $\psi$ also restricts to a diffeomorphism
    $\psi^{I}:(\mathbb{S}_{r}^{2|I|-1},K_{G^{I}})\longrightarrow(\mathbb{S}_{r}^{2|I|-1},K_{F^{I}}),$
    where $K_{G^{I}}$ and $K_{F^{I}}$ are the links defined by the restrictions $F^{I}$ and $G^{I}$, respectively.
2.3. Milnor fibrations
Oka in [26, Theorem 33] proved that strongly non-degenerate convenient mixed functions have Milnor fibrations on the tube and the sphere which are smoothly equivalent. We shall prove an analogous statement for mixed maps under the strong non-degeneracy and linear discriminant conditions. We refer the reader to [9] for further details on the subject of Milnor fibrations.
We recall some definitions. A map germ $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ admits a Milnor fibration on the tube if for each $r>0$ sufficiently small there exists $\delta=\delta(r)>0$ such that the map
(2.2) $F:\mathbb{B}_{r}\cap F^{-1}(\mathbb{S}_{\delta}^{2k-1})\longrightarrow\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r}$
is a locally trivial fibration over its image, where $\mathbb{B}_{r}$ is the open ball centered at the origin with radius $r$ and $\Delta_{r}$ is the discriminant of $F$ restricted to $F(\mathbb{B}_{r})$. For instance, the ICIS condition $\Sigma_{F}\cap F^{-1}(0)\subset\{0\}$ implies the existence of such a fibration by [9, Theorem 2.3]. Hence, the maps in Theorem 2.10 admit a fibration on the tube.
In the notation above, the map $F$ admits a fibration on the sphere if for each sufficiently small $r>0$ there exists a $\delta=\delta(r)>0$ such that the map
(2.3) $\frac{F}{\|F\|}:\mathbb{S}_{r}^{2n-1}\cap F^{-1}(\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r})\longrightarrow\mathbb{S}^{2k-1}_{\delta}\setminus\Delta_{r}$
is a locally trivial fibration over its image.
Lemma 2.13.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a strongly non-degenerate mixed map germ such that for all non-empty $I\subset\{1,\dots,n\}$ the following conditions hold:
  1. (1)
    If $|I|<k$, then $\mathbb{C}^{I}\subset\Sigma_{F}$.
  2. (2)
    If $|I|\geq k$, the coordinate functions satisfy $f^{i,I}\not\equiv 0$.
Then there exists $r_{0}>0$ such that the fibers of the map
$F:\mathbb{B}_{r}\cap F^{-1}\left(\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r}\right)\longrightarrow\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r}$
are transversal to the spheres $\mathbb{S}^{2n-1}_{r}$ for all $0<r<r_{0}$ and $\delta=\delta(r)>0$ sufficiently small.
Proof.
If the statement is false, by [31, Proposition 7] and the Curve Selection Lemma, there exists real analytic curves $w(t)\subset\mathbb{C}^{n}\setminus F^{-1}(\Delta_{F})$ and $\beta_{1}(t),\dots,\beta_{k}(t)\in\mathbb{C}$, $\lambda(t)\subset\mathbb{R}$ non-simultaneously vanishing such that:
$\lambda(t)w(t)=\sum_{j=1}^{k}\beta_{j}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{j}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))+{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{j}{}\mkern 3.0mu\overline{\mkern-3.0muDf^{j}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)).$
Let us denote $I=\{j:w_{j}(t)\not\equiv 0\}$, where $|I|\geq k$, since $w(t)\cap\Sigma_{F}=\emptyset$. This also implies $\lambda(t)\not\equiv 0$. Under these conditions, $f^{i,I}\not\equiv 0$ and we may write the analytic expansions of the coordinate functions and curves:
$f^{i,I}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$ $=c_{i}t^{r_{i}}+o(t),\;c_{i}\in\mathbb{C}^{*}\,,r_{i}\geq 1,$
$\lambda(t)$ $=\lambda_{0}t^{s}+o(t),\;\lambda_{0}\in\mathbb{R}^{*}\,,s\geq 1,$
$w_{i}(t)$ $=a_{i}t^{p_{i}}+o(t),\;a_{i}\in\mathbb{C}^{*}\,,p_{i}\geq 1,$
$\beta_{i}(t)$ $=b_{i}t^{q_{i}}+o(t),\;b_{i}\in\mathbb{C}^{*}\,,q_{i}\geq 1,$
for all $i\in I$. We may suppose $q_{1}+d_{1}\leq\dots\leq q_{k}+d_{k}$, where $d_{i}$ is the weighted homogeneous degree of $f^{i,I}_{P}$. For each $i$, the relation above becomes
$a_{i}\lambda_{0}t^{s+p_{i}}+o(t)=\left(b_{1}\frac{\partial f_{P}^{1,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{1}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f_{P}^{1,I}}{\partial z_{i}}}\right)t^{d_{1}+q_{1}-p_{i}}+o(t)+\dots+\left(b_{k}\frac{\partial f_{P}^{k,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{k}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f_{P}^{k,I}}{\partial z_{i}}}\right)t^{d_{k}+q_{k}-p_{i}}+o(t).$
It follows that:
$\sum_{j=1}^{l}b_{j}\frac{\partial f^{j,I}_{P}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{j,I}_{P}}{\partial z_{i}}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})=\begin{cases}0,&\,\text{if}\;q_{1}+d_{1}-p_{i}-r_{i}<s+p_{i}\\ \lambda_{0}a_{i},&\,\text{if}\;q_{1}+d_{1}-p_{i}-r_{i}=s+p_{i}\end{cases}$
Let us define $K=\{i:q_{1}+d_{1}-p_{i}-r_{i}=s+p_{i}\}$ and we claim that $K\neq\emptyset$. On the contrary, $a\in\mathbb{C}^{*m}$ becomes a critical point of $F_{P}^{I}$ and we apply the first item of Proposition 2.9 to get the first contradiction. The second follows from the following argument. Consider the equations:
(2.4) $\left\langle\sum_{i=1}^{k}\beta_{i}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}F(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))+{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{i}(t){}\mkern 3.0mu\overline{\mkern-3.0muDF}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)),w^{\prime}(t)\right\rangle=\frac{1}{2}\lambda(t)\frac{d}{dt}\|w(t)\|^{2}.$
We may split this dot product as follows:
(2.5) $\left\langle\sum_{i=1}^{k}{}\mkern 3.0mu\overline{\mkern-3.0muD}f^{i},\beta_{i}\frac{dw}{dt}\right\rangle$ $=\sum_{i=1}^{k}\left\langle{}\mkern 3.0mu\overline{\mkern-3.0muD}f^{i}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),b_{i}P\cdot a\right\rangle t^{d_{i}-1+q_{i}}+o(t),$
(2.6) $\left\langle\sum_{i=1}^{k}{}\mkern 3.0mu\overline{\mkern-3.0muDf^{i}},{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{i}\frac{d{}\mkern 3.0mu\overline{\mkern-3.0muw}}{dt}\right\rangle$ $=\sum_{i=1}^{k}\left\langle{}\mkern 3.0mu\overline{\mkern-3.0muDf^{i}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}P\cdot{}\mkern 3.0mu\overline{\mkern-3.0mua}\right\rangle t^{d_{i}-1+q_{i}}+o(t),$
where $P=(p_{1},\dots,p_{n})$ and $a=(a_{1},\dots,a_{n})$. Notice that the Hermitian product satisfies
$\operatorname{Re}\langle v,w\rangle+\operatorname{Re}\langle{}\mkern 3.0mu\overline{\mkern-3.0muv},u\rangle=\operatorname{Re}\langle v,w-{}\mkern 3.0mu\overline{\mkern-3.0muu}\rangle\;\forall\;v,w,u\in\mathbb{C}^{n}.$
Thus, we sum both products (2.5) and (2.6) to obtain that the real part of the coefficient of the lowest degree term of (2.4) is $\lambda_{0}\sum_{j\in K}\|a_{j}\|^{2}p_{j}\neq 0$, provided $K\neq\emptyset$. This allows us to compare the orders:
$d_{1}-1+q_{1}=2p+s-1,$
where $p=\min_{i}\{p_{i}\}$. On the other hand, $s+p=q_{1}+d_{1}-p-r_{l}$ for some index $l\in K$. This leads to $r_{l}=0$, which is a contradiction. ∎
Example 2.14.
Let $f^{1},\dots,f^{k}$ be mixed function germs such that $f^{j}:(\mathbb{C}^{n_{j}},0)\longrightarrow(\mathbb{C},0)$ is strongly non-degenerate, has a critical point at the origin, and $n_{j}>k$ for every $j=1,\dots,k$. Define the mixed map germ $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n_{1}}\times\dots\times\mathbb{C}^{n_{k}},0)\longrightarrow(\mathbb{C}^{k},0)$ on separable variables. We have seen in Example 2.14 that $F$ is strongly non-degenerate. In addition, the matrix $\left(DF\right)^{I}$ does not have maximal rank if $|I|<k$. Moreover, $f^{j,I}\not\equiv 0$ for all $I$ such that $|I|\geq k$. Thus $F$ satisfies the conditions of Lemma 2.13.
Theorem 2.15.
Let $F:\mathbb{C}^{n}\longrightarrow\mathbb{C}^{k}$ be a strongly non-degenerate mixed map as in Lemma 2.13. Suppose further that $F$ has a linear discriminant. Then $F$ admits Milnor fibrations on the tube and the sphere, and these are smoothly equivalent.
Proof.
Lemma 2.13 implies that the map (2.2) is a submersion and the existence of a tube fibration is a consequence of Ehresmann fibration theorem. Moreover, the transversality of the fibers with every small sphere yields that the map (2.3) is a fibration by [9, Proposition 2.12]. Lastly, the equivalence is established in [9, Theorem 2.16]. ∎
3. Contact structures and open books
3.1. Introduction
Let $M$ be a closed orientable odd-dimensional manifold. A contact structure on $M$ is a field $\xi$ of hyperplanes given locally as the kernel $\xi=\operatorname{Ker}(\alpha)$ of a $1$-form $\alpha$ satisfying $\alpha\wedge(d\alpha)^{n}\neq 0$. In other words, $\xi$ is a maximally non-integrable distribution of codimension $1$. The form $\alpha$ is called a contact form. We denote $M$ endowed with this structure by $(M,\xi)$. Moreover, each contact form $\alpha$ is associated with the so-called Reeb vector field $R_{\alpha}$, uniquely determined by the following equations:
$d\alpha(R_{\alpha},-)$ $\equiv 0\,$,
$\alpha(R_{\alpha})$ $\equiv 1\,.$
Let $\rho(z)=\sum_{i=1}^{n}\|z_{i}\|^{2}$ be the square distance function. The spheres $\rho^{-1}(r^{2})=\mathbb{S}_{r}^{2n-1}$ are endowed with a contact structure called natural or canonical, denoted by $\xi_{r}$, and associated with the restriction of the following contact form:
(3.1) $\alpha$ $=2\sum_{i=1}^{n}(x_{i}dy_{i}-y_{i}dx_{i})=-{\bf i}\sum_{i=1}^{n}({}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}dz_{j}-z_{j}d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}),$
where we take coordinates $z_{i}=(x_{i},y_{i})$ of $\mathbb{C}^{n}$. For each $p\in\mathbb{S}_{r}^{2n-1}$, the subspace $\xi_{r}(p)$ corresponds to the subspace of $T_{p}\mathbb{S}_{r}^{2n-1}$ invariant by the complex structure $J$, where $J^{2}=-\operatorname{Id}$. The associated Reeb vector field is
(3.2) $R=\frac{1}{2r^{2}}\sum_{j=1}^{n}\left(x_{j}\frac{\partial}{\partial y_{j}}-y_{j}\frac{\partial}{\partial x_{j}}\right)=\frac{1}{2\rho}\sum_{j=1}^{n}\left(z_{j}\frac{\partial}{\partial z_{j}}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right).$
Let $(V,0)\subset\mathbb{C}^{n}$ be a complex isolated singularity germ at the origin. The restriction of the square distance function to the complex manifold $V\setminus\{0\}$ induces a contact structure on the links $K_{r}=V\cap\mathbb{S}^{2n-1}_{r}$ so that $K_{r}$ is a contact submanifold of $\mathbb{S}^{2n-1}_{r}$ for each sufficiently small $r>0$. We refer the reader to [6] and [5] for details.
Remark 3.1 (Orientations).
We state the following convention. On the spheres $\mathbb{S}_{r}^{2n-1}$ and the links admitting a contact structure, the positive orientation is that given by the volume form $\lambda$ in (3.1).
Recall that two contact manifolds $(M_{1},\xi_{1})$ and $(M_{2},\xi_{2})$ are contactomorphic, or isomorphic, if there exists a diffeomorphism $\phi:M_{1}\longrightarrow M_{2}$ such that $d\phi(\xi_{1})=\xi_{2}$. Varchenko in [37] showed that the isotopy type of the contact manifold $K_{V}$ constructed from a complex isolated singularity germ $V$ does not depend on the embedding and the radius $r$ of the sphere given by the strictly plurisubharmonic function. Henceforth we shall denote the link of a variety $V$ or map germ $F$ with an isolated singularity at the origin by $K_{V}$ or $K_{F}$, respectively. An oriented contact manifold contactomorphic to such a holomorphic link is called Milnor fillable. This name is a reference to the fact that a complex link endowed with the natural contact structure is the boundary of the Milnor fiber with its natural symplectic structure. For details, see [30, Section 6].
An open book on an oriented manifold $M$ is a pair $(N,\theta)$ such that $N\subset M$ is a codimension 2 orientable submanifold with trivial normal bundle and $\theta:M\setminus N\longrightarrow\mathbb{S}^{1}$ is a locally trivial fibration which coincides with the angular coordinate on a trivial tubular neighborhood of $N$. We suppose that $N$ has the boundary orientation induced by the fibers of $\theta$. Open books are closely related to contact manifolds as proved by Giroux in [11].
Definition 3.2.
Let $(M,\xi)$ be an oriented closed manifold supporting a contact structure $\xi$ defined by a 1-form $\alpha$. We say that $\xi$ is adapted to, or carried by, an open book $(N,\theta)$ if:
  1. (1)
    The restriction of $\alpha$ to $N$ is a positive contact form.
  2. (2)
    The 2-form $d\alpha$ defines a symplectic form on each fiber of $\theta$.
Lemma 3.3 (Lemma 2.2, [6]).
Let $M$ be a closed oriented manifold and $\psi:M\longrightarrow\mathbb{C}$ a differentiable function. Let $\Theta_{\psi}:=\psi/\|\psi\|:M\setminus\psi^{-1}(0)\longrightarrow\mathbb{S}^{1}$ and suppose there exists $\eta>0$ such that:
  1. (1)
    $d\left(\Theta_{\psi}\right)\neq 0$ if $\|\psi\|\geq\eta$, and
  2. (2)
    $d\psi\neq 0$ if $\|\psi\|\leq\eta$.
Then $(\psi^{-1}(0),\Theta_{\psi})$ is an open book in $M$.
For instance, on a 3-dimensional closed oriented manifold, any contact structure is carried by some open book. Moreover, two positive contact structures carried by the same open book are isotopic. Additionally, any Milnor fillable oriented 3-manifold admits a unique Milnor fillable contact structure up to contactomorphism. Also, in dimension 3, contact structures are divided into two types: overtwisted and tight. For instance, Milnor fillable manifolds and spheres endowed with the natural structure are the first examples of tight structures. For details, see [30, Theorem 5.21 and Section 6]. A classification of overtwisted structures is developed in [10].
Let $(V,0)\subset\mathbb{C}^{n}$ be a complex germ set with an isolated singularity at the origin. Let $K_{V}=V\cap\mathbb{S}_{r}^{2n-1}$ be the link, where $r>0$ is sufficiently small. For any holomorphic function germ $h:(V,0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin, one can consider the argument function $\Theta_{h}:=h/\|h\|:K_{V}\setminus h^{-1}(0)\longrightarrow\mathbb{S}^{1}$. The authors showed in [6, Theorem 3.9] that it is adapted to the natural contact structure on $K_{V}$. We remark that the underlying constructions in this theorem depend on the holomorphic setting and one cannot expect an analogous statement for the real case (see [6, Lemmas 3.6 and 3.7]). Therefore, additional hypotheses are needed to obtain open books on mixed ICIS (see Subsection 3.3).
3.2. Links of mixed ICIS
In this section, we extend for mixed maps the constructions performed for mixed functions in [28] regarding contact structures. Let $f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ be a mixed function. We consider the following notation:
$\frac{\partial f}{\partial z_{i}}=f_{z_{i}}\,,\;\;\frac{\partial f}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}=f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}.$
We begin with a lemma used later for some computations.
Lemma 3.4 (Section 3.3, [28]).
Let $\rho$ and $\alpha$ be as in (3.1) and $f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=g(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})+{\bf i}h(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ be a mixed function.
  1. (1)
    The 2-form $d\rho\wedge\alpha$ is given by:
    $d\rho\wedge\alpha={\bf i}\sum_{i,j}A_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j},$
    where $A_{i,j}=2{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}z_{j}$.
  2. (2)
    The 2-form $dg\wedge dh$ is given by:
    $dg\wedge dh={\bf i}\sum_{i,j}B_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}+R,$
    where $R$ is a linear combination of other types of 2-forms and
    $B_{i,j}=\frac{1}{2}\left(f_{z_{i}}\overline{f_{z_{j}}}-\overline{f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right).$
  3. (3)
    The 4-form $d\rho\wedge\alpha\wedge dg\wedge dh$ is given by
    $d\rho\wedge\alpha\wedge dg\wedge dh=-\sum_{i,j}C_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\wedge dz_{j}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}+S,$
    where $S$ is a linear combination of other types of 4-forms and
    $C_{i,j}=\|{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}f_{z_{j}}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{b}f_{z_{i}}\|^{2}-\|z_{i}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}-z_{j}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\|^{2}.$
  4. (4)
    One has the following equality:
    $d\rho\wedge\alpha\wedge d\alpha^{n-2}\wedge dg\wedge dh(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ $=\kappa(n)C(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
    where $\kappa(n)={\bf i}^{n}2^{n-2}(n-2)!$ and $C(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{1\leq i<j\leq n}C_{i,j}$.
Lemma 3.5.
Let $f^{1},\dots,f^{k}:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ be mixed function germs. The following equality verifies:
$d\rho\wedge\alpha\wedge df_{1}^{1}\wedge df_{2}^{1}\wedge\dots\wedge df_{1}^{k}\wedge df_{2}^{k}=\sum_{j_{1},\dots,j_{k}}D_{j_{1},\dots,j_{k}}dz_{j_{1}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}\wedge\dots\wedge dz_{j_{k}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{k}}+T,$
where $T$ is a linear combination of other types of $2k+2$-forms and
$D_{j_{1},\dots,j_{k}}$ $=2^{-k+2}\cdot C_{j_{1},j_{2}}B^{1}_{j_{3},j_{3}}\dots B^{k-1}_{j_{k},j_{k}}$
$=\left(\|{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}f^{1}_{z_{j_{2}}}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{2}}f^{1}_{z_{j_{1}}}\|^{2}-\|z_{j_{1}}f^{1}_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{2}}}-z_{j_{2}}f^{1}_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}}\|^{2}\right)\left(\|f^{2}_{z_{j_{3}}}\|^{2}-\|f^{2}_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{3}}}\|^{2}\right)\dots\left(\|f^{k-1}_{z_{j_{k}}}\|^{2}-\|f^{k-1}_{z_{j_{k}}}\|^{2}\right).$
Proof.
We operate the wedge product of the form $d\rho\wedge\alpha\wedge df_{1}^{1}\wedge df_{2}^{1}$ with the forms $df_{1}^{m}\wedge df_{2}^{m}$ putting together the terms $dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}$. Moreover, the form $T$ is the result of the wedge product of $S$ and terms with the form $B_{i,j}^{m}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}$, where $B_{i,j}^{m}$ is the expression associated to the coordinate function $f^{m}$. ∎
Let us denote the sum of $D_{j_{1},\dots,j_{k}}$ with $j_{1},\dots,j_{k}$ distinct by $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$. Notice that if $k=1$, it is nothing but the sum $C(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ in item 4 of Lemma 3.4.
Corollary 3.6.
One has the following expression:
$d\rho\wedge\alpha\wedge d\alpha^{n-(k+1)}\wedge df_{1}^{1}\wedge df_{2}^{1}\wedge\dots\wedge df_{1}^{k}\wedge df_{2}^{k}=\kappa(n)D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
where $\kappa(n)$ is a constant depending only on $n$.
Proof.
We have that $d\alpha=-2i\sum_{s=1}^{n}dz_{s}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{s}$ and so
$(d\alpha)^{n-(k+1)}=\sum_{s_{1},\dots,s_{n-(k+1)}}dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge d\widehat{z}_{s_{1}}\wedge d\widehat{{}\mkern 3.0mu\overline{\mkern-3.0muz}}_{s_{1}}\wedge\dots\wedge d\widehat{z}_{s_{n-(k+1)}}d\widehat{{}\mkern 3.0mu\overline{\mkern-3.0muz}}_{s_{n-(k+1)}}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
where $\widehat{z}_{s_{j}}$ are removed variables. Notice that the permutation of a pair $dz_{s_{j}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{s_{j}}$ does not change the sign of the form. Moreover, there are (n(k+1))!(n-(k+1))! terms in the sum above and we claim that $(d\alpha)^{n-(k+1)}\wedge T=0$. Indeed, a form that appears in the sum $T$ must involve at least $k+1$ indices and, applying the wedge product with the terms of $(d\alpha)^{n-(k+1)}$, we always find repetitions and the assertion follows. ∎
Likewise in [28, Section 3.4], we have the following definition.
Definition 3.7.
A mixed map germ $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ is called holomorphic-like (respectively anti-holomorphic-like) if $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\geq 0$ (respectively $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\leq 0$) for a sufficiently small neighborhood of the origin. If the inequalities are strict, then we call $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ strictly (anti-)holomorphic-like.
Lemma 3.8.
Let $F=(f^{1},\dots,f^{k}):\mathbb{C}^{n}\longrightarrow\mathbb{C}^{2}$ be a holomorphic map germ and $\phi_{a,b}$ a homogeneous mixed covering, where $a>b$ (respectively, $a<b)$. Then the mixed map $G(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\phi_{a,b}^{*}F$ is holomorphic-like (respectively, anti-holomorphic).
Proof.
Let us denote $\phi:=\phi_{a,b}$. For each $j$, the following equations hold:
$g^{l}_{w_{j}}$ $=f^{l}_{w_{j}}\left(\phi(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})\right)aw_{j}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{b},$
$g^{l}_{{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}}$ $=f^{l}_{w_{j}}\left(\phi(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})\right)bw_{j}^{a}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{b-1}.$
Moreover:
$\|g^{l}_{w_{j}}\|^{2}-\|g^{l}_{{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}}\|^{2}=(a^{2}-b^{2})\|\phi^{*}f^{l}_{w_{j}}\|^{2}\|w_{j}\|^{2(a+b-1)}.$
Substituting this in (3.5), the real function $D_{j_{1},\dots,j_{k}}$ is equal to
$(a^{2}-b^{2})^{k}\|w_{j_{3}}\dots w_{j_{k}}\|^{2(a+b-1)}\|w_{j_{1}}w_{j_{2}}\|^{2}\|w_{j_{1}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{1}}^{b-1}\phi^{*}f_{z_{j_{1}}}-w_{j_{2}}^{a-1}w_{j_{2}}^{b-1}\phi^{*}f_{z_{j_{2}}}\|^{2}\|\phi^{*}f^{2}_{z_{j_{3}}}\|^{2}\dots\|\phi^{*}f^{k-1}_{z_{j_{k}}}\|^{2}.$
Theorem 3.9.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a partially non-degenerate holomorphic map germ as in Theorem 2.10 and $\phi_{a,b}$ a homogeneous mixed covering with $a>b$ (respectively, $a<b)$. Then the link $K_{G}$ of the mixed ICIS given by $G=\phi_{a,b}^{*}F$ is a positive (respectively, negative) contact submanifold of the sphere for every sufficiently small $r>0$.
Proof.
We follow the proof of [28, Theorem 1]. We consider the case $a>b$ and the other is analogous. Theorem 2.10 implies that the link $K_{G}$ is a real smooth manifold of codimension $2k+1$ for sufficiently small $r>0$. Let us denote $G=(g^{1},\dots,g^{k})$ and $g_{1}^{i},g_{2}^{i}$ the real and imaginary parts of $g^{i}$, respectively. Notice that $K_{G}$ is a complete intersection defined by $g_{1}^{i}=g_{2}^{i}=0$ and $\rho-r^{2}=0$, for $i=1,\dots,k$. This implies that there exists a local coordinate system formed by $g_{1}^{i},g_{2}^{i},\rho$, and other real analytic functions $h_{2k+2},\dots,h_{2n}$. Therefore, the condition $\alpha\wedge(d\alpha)^{n-(k+1)}\not\equiv 0$ is equivalent to
$d\rho\wedge\alpha\wedge d\alpha^{n-(k+1)}\wedge dg_{1}^{1}\wedge dg_{2}^{1}\wedge\dots\wedge dg_{1}^{k}\wedge dg_{2}^{k}\not\equiv 0.$
[…]
$\begin{cases}&\|w_{j_{3}}\dots w_{j_{k}}\|^{2(a+b-1)}\|w_{j_{1}}w_{j_{2}}\|^{2}\|w_{j_{1}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{1}}^{b-1}\phi^{*}f_{z_{j_{1}}}-w_{j_{2}}^{a-1}w_{j_{2}}^{b-1}\phi^{*}f_{z_{j_{2}}}\|^{2}\|\phi^{*}f^{2}_{z_{j_{3}}}\|^{2}\dots\|\phi^{*}f^{k-1}_{z_{j_{k}}}\|^{2}\equiv 0\\ &\phi^{*}f^{j}\equiv 0\;\,\text{for all}\;j=1,\dots,k,\end{cases}$
where $f^{i}$ and $f_{z_{j_{l}}}^{i}$ are restricted to the curve $w(t)$. Let $I=\{j:w_{j}(t)\not\equiv 0\}$. By the non-degeneracy condition of the coordinate functions restricted to $V_{G}$ and Lemma 2.2, there exist $j_{1},\dots,j_{k}\in I$ such that $f^{l}_{z_{j_{k}}}(w(t))\not\equiv 0$ for each $l=1,\dots,k$. Then $D_{j_{1},\dots,j_{k}}\equiv 0$ and
(3.3) $\left\|w_{j_{1}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{1}}^{b-1}\phi^{*}f^{1}_{z_{j_{1}}}-w_{j_{2}}^{a-1}w_{j_{2}}^{b-1}\phi^{*}f^{1}_{z_{j_{2}}}\right\|^{2}\equiv 0.$
Let $J=\{j_{l}:w_{j_{l}}(t)\not\equiv 0\}$ and define the curve $\gamma(t)=w_{j_{l}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b-1}\phi^{*}f_{z_{j_{l}}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$. Note that $\gamma(t)\not\equiv 0$ and any $j_{l}\in J$ such that $f_{z_{j_{l}}}\left(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\right)\not\equiv 0$ defines the same curve by the equation (3.3). Let $v_{j_{l}}(t)=w_{j_{l}}^{\prime}(t)$. We have that:
$0$ $=\frac{dg(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))}{dt}$
$=\sum_{l}^{n}g_{w_{j_{l}}}\frac{dw_{j_{l}}}{dt}+g_{{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}}\frac{d{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}}{dt}$
$=\sum_{l}^{n}f_{z_{j_{l}}}(\phi(w(t),w(t)))\left(aw_{j_{l}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b}v_{j_{l}}+bw_{j_{l}}^{a}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b-1}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}\right)$
$=\sum_{l}^{n}f_{z_{j_{l}}}(\phi)w_{j_{l}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b-1}\left(a{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j}+bw_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}\right)$
$=\gamma(t)\sum_{l}^{n}\left(a{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}+bw_{j}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}\right).$
Notice that the sum above can be taken on $j_{l}\in J$, because if $j_{l}\notin J$ we have null terms. Since this last term is zero, summing up $\sum_{l}aw_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}$ and $\sum_{j_{l}}b{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}$ does not change the equality. Thus
$0=\gamma(t)\left(\frac{a+b}{2}\right)\sum_{l}^{n}w_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}+{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}.$
Taking the derivative of $\|w(t)\|^{2}=\sum_{l}^{n}w_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}$ we obtain:
$\frac{d\|w(t)\|^{2}}{dt}=\sum_{l}^{n}w_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}+{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}=0.$
Since $\lim_{t\to 0}w(t)=0$, it implies $w(t)\equiv 0$, which is a contradiction and the result follows. ∎
Remark 3.10.
  1. (1)
    In [28, Theorem 1], the mixed function is the pullback of non-degenerate convenient holomorphic functions. As we have seen, these two conditions together are generalized by partial non-degeneracy, which is also sufficient to obtain the same result in the case $k=1$.
  2. (2)
    The first paragraph in the previous proof implies that a mixed link determined by a mixed ICIS is a positive (respectively, negative) contact submanifold of the sphere $\mathbb{S}_{r}^{2n-1}\subset\mathbb{C}^{n}$ if and only if the associated mixed map is (resepctively, anti-)holomorphic-like.
  3. (3)
    We shall refer to the contact structure above as natural and denote it by $\xi_{r}$.
A direct consequence of the Curve Selection Lemma implies that the contact type of the link does not depend, up to a contactomorphism, on the holomorphic coordinate system. More generally, we can formulate the statement as follows.
Proposition 3.11.
Let $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ be a strictly (anti-)holomorphic-like mixed map and $\varphi:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ a real analytic diffeomorphism. Then $H(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\varphi^{*}G$ is also a strictly (anti-)holomorphic-like mixed map.
Proof.
Let us denote $D_{H}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=D\circ\varphi(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ the associated real function of $H$ and suppose the statement is false. Moreover, let us introduce coordinates $\varphi(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$. By the Curve Selection Lemma, there exists an open neighborhood $U$ of the origin and a real analytic curve $\lambda:(0,1]\longrightarrow U\subset\mathbb{C}^{n}$ such that:
$\begin{cases}&\lim_{t\to 0}\lambda(t)=0\Longrightarrow\lim_{t\to 0}\varphi(\lambda(t))=0,\\ &H(z(t),{}\mkern 3.0mu\overline{\mkern-3.0muz}(t))\equiv 0\Longrightarrow G(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0,\\ &D_{H}(z(t),{}\mkern 3.0mu\overline{\mkern-3.0muz}(t))\equiv 0\Longrightarrow D(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0,\end{cases}$
where $z(t),w(t)$ denote the restrictions of the coordinates to the curve $\lambda(t)$. Then $\varphi(\lambda(t))$ is a real analytic curve on $V_{G}$ on which the link $K_{G}=V_{G}\cap\mathbb{S}_{r}^{2n-1}$ of $g$ is not a contact submanifold for all $r>0$ such that $\mathbb{S}_{r}^{2n-1}\subset U$. But this is a contradiction with the initial assumption. ∎
Remark 3.12.
Notice that the existence of the natural contact structure on a link defined by a (anti-)holomorphic-like mixed map only makes sense if the ambient space has a holomorphic structure.
This proposition allows us the following conclusion. Let $(V,x)\subset(M,m)$ be the germ of a mixed variety with an isolated singularity at $x$ and $(M,m)$ the germ of a complex analytic manifold. If there exists a coordinate system $z$ of $M$ for which $K_{V}=V\cap\mathbb{S}_{r,z}^{2n-1}$ is a contact manifold, where $\mathbb{S}_{r,z}^{2n-1}$ is the usual sphere on the coordinate system $z$, then the same assertion holds for any other holomorphic local coordinate system of $M$.
3.3. Open books
We prove in this section the existence of open books adapted to the natural contact structures on mixed links of ICIS. Given a mixed function $g$ with an isolated singularity at the origin, we establish an extra condition related to the mixed ICIS $G=\phi^{*}F$ that allows us to derive an open book that is further adapted. This hypothesis is based on the proof of [5, Proposition 3.2], in the holomorphic non-isolated singularity context, and [28, Theorem 4].
Firstly, we recall the main steps of a mixed-version construction of [6] developed in [28]. Let $\alpha$ be the natural contact form (3.1) and $g:\mathbb{C}^{n}\longrightarrow\mathbb{C}$ be a mixed function. We modify $\alpha$ by
(3.4) $\alpha_{c}=e^{-c\|g\|^{2}}\cdot\alpha,$
where $c>0$. Notice that the corresponding hyperplane field $\xi$ is not modified. Let $\pi^{\perp}:\mathbb{C}^{n}\longrightarrow\mathbb{C}\cdot R$ be the projection on the line generated by the Reeb vector field and the orthogonal complement $\pi(v)=v-\pi^{\perp}(v)$. Recall the gradient vector fields (1.1):
$Dg$ $=\left(\frac{\partial g}{\partial w_{1}},\dots,\frac{\partial g}{\partial w_{1}}\right),$
${}\mkern 3.0mu\overline{\mkern-3.0muD}g$ $=\left(\frac{\partial g}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muw}_{1}},\dots,\frac{\partial g}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muw}_{1}}\right).$
Write
$g{}\mkern 3.0mu\overline{\mkern-3.0muD}g$ $=\pi(g{}\mkern 3.0mu\overline{\mkern-3.0muD}g)+\pi^{\perp}(g{}\mkern 3.0mu\overline{\mkern-3.0muD}g),$
${}\mkern 3.0mu\overline{\mkern-3.0mug}{}\mkern 3.0mu\overline{\mkern-3.0muD}g$ $=\pi({}\mkern 3.0mu\overline{\mkern-3.0mug}{}\mkern 3.0mu\overline{\mkern-3.0muD}g)+\pi^{\perp}({}\mkern 3.0mu\overline{\mkern-3.0mug}{}\mkern 3.0mu\overline{\mkern-3.0muD}g).$
Let
$v_{1}=\pi(gDg),\;\;\;\;v_{2}=\pi({}\mkern 3.0mu\overline{\mkern-3.0mug}{}\mkern 3.0mu\overline{\mkern-3.0muD}g).$
By [28, p.266], the expression of the Reeb vector field $R_{c}$ of $\alpha_{c}$ becomes:
(3.5) $\|g\|^{2}d\Theta_{g}(R_{c})=e^{c\|g\|^{2}}\|g\|^{2}d\Theta_{g}(R)+\frac{ce^{c\|g\|^{2}}}{2}\left(\|v_{1}\|^{2}-\|v_{2}\|^{2}\right).$
Suppose that $g=\phi^{*}f$, where $\phi$ is a homogeneous mixed covering and $f$ a holomorphic function. By [28, Lemma 4], $\|v_{1}\|^{2}\geq\|v_{2}\|^{2}$ and the equality holds if and only if $\nabla\Theta_{g}=\lambda R$ for some $\lambda\in\mathbb{C}$, where
$\nabla\Theta_{g}=i\left(\frac{{}\mkern 3.0mu\overline{\mkern-3.0mug}_{z_{1}}}{{}\mkern 3.0mu\overline{\mkern-3.0mug}}-\frac{g_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}}{g},\dots,\frac{{}\mkern 3.0mu\overline{\mkern-3.0mug}_{z_{n}}}{{}\mkern 3.0mu\overline{\mkern-3.0mug}}-\frac{g_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n}}}{g}\right).$
Theorem 3.13.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ and $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ be partially non-degenerate holomorphic map and function germs, respectively. Let $\phi_{a,b}$ and $\phi_{c,d}$ be homogeneous mixed coverings, where $a>b$ and $c>d$, and define the pullbacks $G(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\phi_{a,b}^{*}F$ and $g(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\phi_{c,d}^{*}f$. Suppose further that $g(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$ defines with $G(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$ a mixed ICIS germ $\Psi:=(G,g):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k+1},0)$. Then the restriction
$\Theta_{g}:=g/\|g\|:K_{G}\setminus K_{g}\longrightarrow\mathbb{S}^{1}$
of the argument of $g$ to the link $K_{G}=V_{G}\cap\mathbb{S}^{2n-1}_{r}$ defines an open book adapted to the natural contact structure, where $r>0$ is sufficiently small.
Proof.
First, since the map $\Psi=(G,g):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k+1},0)$ is an isolated complete intersection singularity, there exist $r_{0}>0$ and $\eta>0$ such that $\Psi^{-1}(s,t)$ intersect the sphere $\mathbb{S}^{2n-1}_{r}$ transversely for all $r<r_{0}$ and $\|(s,t)\|<\eta$. Whence, the fibers $g^{-1}(t)$ intersect $K_{G}$ transversely for $t$ sufficiently small. Recall that $g$ has an isolated singularity at the origin by Proposition 2.3. By Lemma 3.3, this implies that $\Theta_{g}$ defines an open book in $K_{G}$. On the other hand, by Theorem 3.9, the link $K_{g}=V_{g}\cap\mathbb{S}^{2n-1}_{r}$ is a contact submanifold as well as its restriction to $K_{G}$. Considering our convention for the orientations, it remains to verify that the fibers of $\Theta_{g}$ have the natural symplectic structure. We shall apply the same strategy of [6, Theorem 3.9] and [28, Theorem 4]. That is, we consider the modification $\alpha_{c}$ in (3.4) which induces the same hyperplane distribution but satisfies $d\Theta(R_{c})>0$. Define
$Z_{\delta}=\{w\in K_{G}\setminus N_{\delta}:d\Theta_{g}(R)\leq 0\},$
where $N_{\delta}\subset K_{G}$ is a tubular neighborhood of $K_{g}$ in $K_{G}$. The regularity of $\Theta_{g}$ implies it is a normal angular coordinate on $N_{\delta}$. Recall equation (3.5). We shall see that one of the following conditions holds:
  1. (1)
    $\|v_{1}\|>\|v_{2}\|$; or
  2. (2)
    $d\Theta_{g}(R)>0$ when $\|v_{1}\|=\|v_{2}\|$.
For the first case, it is enough to choose a sufficiently large $c>0$ to make $d\Theta_{g}(R_{c})>0$. Moreover, we have claimed that $\|v_{1}\|=\|v_{2}\|$ if and only if $\nabla\Theta_{g}(w)=\lambda R(w)$ and, in this case, $d\Theta_{g}(R)=\operatorname{Re}\lambda\|R\|^{2}$. By [28, Lemma 5], if $w\in Z_{\delta}$ is a solution for this equation, then $\operatorname{Re}\lambda>0$. We conclude that $d\Theta_{g}(R)>0$. ∎
4. Natural and Milnor fillable structures
Some classes of mixed maps are related by topological and smooth equivalences with holomorphic maps, as for instance the mixed Hamm ICIS in Subsection 2.2. This implies the existence of a contact structure induced from that in the complex link, which is Milnor fillable. If they are further endowed with the natural contact structure as a mixed singularity, when this exists, we address the problem of comparing them. We prove that in the case of the mixed Hamm ICIS, these are isotopic.
Let $G:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed ICIS germ and $(V,0)\subset\mathbb{C}^{n}$ be a complex germ with an isolated singularity at the origin. Let $K_{G}$ be the mixed link and suppose the existence of a map germ $\varphi:(\mathbb{C}^{n},K_{V})\longrightarrow(\mathbb{C}^{n},K_{G})$ which is a diffeomorphism on $K_{V}$. One can define a contact structure on $K_{G}$ by setting $\xi_{G}=d\phi(\xi_{V})$. This occurs for mixed Hamm ICIS in Theorem 2.12. See also [26] and [15].
If there exists another diffeomorphism $\psi:(\mathbb{C}^{n},K_{V})\longrightarrow(\mathbb{C}^{n},K_{G})$ and we set a contact structure $\xi_{G}^{\prime}=d\psi(\xi_{V})$ on $K_{G}$ induced by $\psi$, it is clear that $\xi_{G}^{\prime}$ and $\xi_{G}$ are contactomorphic. Moreover, if $(\Theta,N)$ is an open book adapted to $\xi_{V}$, the induced one is $\psi^{*}\Theta$, where $\psi=\varphi^{-1}$ and the binding is $\phi(N)$. Recall that we have set $\alpha$ as the contact form (3.1) on the sphere.
Proposition 4.1.
There exists an open book adapted to the induced Milnor fillable contact structure for which $d\alpha$ defines a symplectic form on each fiber.
Proof.
Considering the notation above, we must show that $d(\psi^{*}\Theta)(R)$ is non-vanishing, where $R$ is the Reeb vector field (3.2). More precisely, by [6, Theorem 3.9], it is enough to find a holomorphic function germ $h$ with an isolated singularity such that the above condition is satisfied, with $\Theta=h/\|h\|$. Recall that
$R$ $=\sum_{j}z_{j}\frac{\partial}{\partial z_{j}}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}},$
$d\Theta$ $=\frac{\partial h}{h}-\frac{{}\mkern 3.0mu\overline{\mkern-3.0mu\partial}{}\mkern 3.0mu\overline{\mkern-3.0muh}}{{}\mkern 3.0mu\overline{\mkern-3.0muh}},$
since $\mkern 3.0mu\overline{\mkern-3.0mu\partial}h=\partial{}\mkern 3.0mu\overline{\mkern-3.0muh}=0$ because $h$ is holomorphic. We obtain that
$d(\Theta\circ\psi)_{z}(R)=\sum_{i,j}$ $\frac{\partial h}{\partial z_{j}}\frac{z_{j}}{h}\left[\frac{\partial}{\partial z_{j}}\left(\frac{\psi_{i}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2}\right)-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\left(\frac{\psi_{i}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2}\right)\right]$
- $\frac{\partial{}\mkern 3.0mu\overline{\mkern-3.0muh}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\frac{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}{{}\mkern 3.0mu\overline{\mkern-3.0muh}}\left[\frac{\partial}{\partial z_{j}}\left(\frac{\psi_{i}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2i}\right)-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\left(\frac{\psi_{i}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2i}\right)\right].$
If $\alpha_{j}=\frac{\partial h}{\partial z_{j}}\frac{z_{j}}{h}$, we may rewrite the expression as
(4.1) $\frac{{\bf i}}{2}d(\Theta\circ\psi)_{z}(R)$ $=\sum_{i,j}\alpha_{j}\left(\frac{\partial\operatorname{Re}(\psi_{i})}{\partial z_{j}}-\frac{\partial\operatorname{Re}(\psi_{i})}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right)-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{j}\left(\frac{\partial\operatorname{Im}(\psi_{i})}{\partial z_{j}}-\frac{\partial\operatorname{Im}(\psi_{i})}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right)$
$=\sum_{i,j}\alpha_{j}\frac{\partial\operatorname{Re}(\psi_{i})}{\partial y_{j}}-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{j}\frac{\partial\operatorname{Im}(\psi_{i})}{\partial y_{j}}.$
Let $h=z_{1}$ and set $r_{1}=\sum_{i}\frac{\partial\operatorname{Re}\psi_{i}}{\partial y_{1}}$ and $r_{2}=\sum_{i}\frac{\partial\operatorname{Im}\psi_{i}}{\partial y_{1}}$. By the local form of immersions, we may suppose the Jacobian matrix of $\psi$ in the coordinates $(x,y)$ has the following local form:
$\begin{pmatrix}(D\psi)_{2n-(2k+1)\times 2n-(2k+1)}&{0}_{2n-(2k+1)\times 2k+1}\\ {0}_{(2k+1)\times 2n-(2k+1)}&{0}_{2k+1\times 2k+1}\end{pmatrix}$,
where $D\psi$ is an invertible square matrix. If (4.1) is zero at a point $p$ and $\Theta$ is defined by $h$, then $r_{1}-r_{2}=0$. Multiplying the rows corresponding to the imaginary parts and summing up, we obtain a null column and this leads to a contradiction with the fact that $\psi$ is a diffeomorphism on $K_{V}$. ∎
For the next results, we suppose the link of the mixed map $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is ambient Milnor fillable.
Corollary 4.2.
Suppose that $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is strictly holomorphic-like. If the restriction of $\xi_{r}$ to the binding $\phi(N)$ is a contact structure, then the induced Milnor fillable and the natural contact structures are isotopic.
Proof.
In this case, $\xi_{r}$ and $\xi_{G}$ are both adapted to $(\psi^{*}(\Theta),\phi(N))$. Consider a fiber of $\psi^{*}\Theta$ and note that it is endowed with two symplectomorphic structures, namely, $\psi^{*}(d\alpha)$ and $d\alpha$. Besides being isotopic, the symplectic structures on the completions are also symplectomorphic, and thus the result follows from [11, Proposition 9]. ∎
In the case nk=2n-k=2, the binding $\phi(N)$ has dimension $1$ and we conclude the following.
Corollary 4.3.
Suppose that $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is strictly holomorphic-like and nk=2n-k=2. Then the induced Milnor fillable and the natural contact structures are isotopic.
Theorem 4.4.
Let $G:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed Hamm ICIS germ given as the pullback $G=\phi_{a,b}^{*}H$ of some complex map germ $H:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ by a homogeneous mixed covering $\phi_{a,b}$, where $a>b$. Then the induced Milnor fillable structure is isotopic with the natural contact structure.
Proof.
The Milnor fillable contact structure is induced from the diffeomorphism $\psi:=\psi_{1}$ of Theorem 2.12 and the binding $N=K_{F}\cap\{z_{1}=0\}$ is mapped to $\psi(N)=K_{G}\cap\{z_{1}=0\}$, where $F$ is the associated Hamm ICIS. But $\psi(N)$ is the link of $G^{I}$, where $I=\{2,\dots,n\}$. Furthermore, $G^{I}$ is a strictly holomorphic-like mixed function and so $\psi(N)$ is a contact submanifold of $\mathbb{S}_{r}^{2n-3}$. The result follows from Corollary 4.2. ∎
Acknowledgments
We thank Patrick Popescu-Pampu for useful discussions and for calling to our attention the fact remarked in Proposition 3.11. We also thank Maria Aparecida Ruas, Cidinha, for helpful conversations.
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data Availability
Data availability is not applicable to this article.